# Evidence for Inflation

### v1

@@ Abstract
Since our constraining results have revealed that the GUT scale can be reached during inflation within $2\,\sigma$ level, the particle production of dark matter around the GUT scale is allowed by observations. As a consequence, our constraints could inspire the exploration of dark matter model building at extremely high energy scales, and intriguingly, provide a new window to investigate the physics before inflation.
@@
Acknowledgements. DW is supported by the CDEIGENT Fellowship of Consejo Superior de Investigaciones Científicas (CSIC). DW warmly thanks the usage of Planck Planck:2018vyg , WMAP WMAP:2012nax , ACT ACT:2020gnv , SPT SPT-3G:2022hvq , BK18 BICEP:2021xfz and DESI DESI:2024mwx data.
S1 Supplementary Material
S1.1 A. $5\,\sigma$ contours from BK18
In Fig.S1, we present the constrained ($n_{s}$, $\log_{10}r$) parameter space from the BK18 observations within the $5\,\sigma$ confidence level (CL). It is easy to see that PGWs are detected at beyond the $5\,\sigma$ CL. This can act as an important evidence of PGWs independent of the four CMB experiments including Planck, WMAP, ACT and SPT.
S1.2 B. Inflationary equation of state
In Fig.S2, we show the one-dimensional posterior distribution of inflationary equation of state (EoS). It is easy to see that the nature of inflation is quintessence, i.e., $\omega>-1$, at beyond the $5\,\sigma$ CL.
S1.3 C. CMB B-mode polarization maps
In Fig.S3, we show the CMB B-mode polarization maps for the best fit, $1\,\sigma$, $2\,\sigma$ and $3\,\sigma$ upper bounds of $\log_{10}r$ from the Planck-only constraint.
Figure S1: The two-dimensional marginalized posterior distributions of the parameter pair ($n_{s}$, $\log_{10}r$) from the BK18 observations within the $5\sigma$ confidence level.
Figure S2: The one-dimensional marginalized posterior distribution of inflationary equation of state from the Planck data.
Figure S3: The CMB B-mode polarization maps for the best fit (upper left), $1\,\sigma$ (upper right), $2\,\sigma$ (lower left) and $3\,\sigma$ (lower right) upper bounds of $\log_{10}r$ from the Planck-only constraint.
Figure S4: The one-dimensional marginalized posterior distribution of the logarithmic tensor-to-scalar ratio ${\rm{log}_{10}}r$ from the IPTA DR2.
S1.4 D. The inconsistency of constraints on $r$ between CMB and pulsar timing arrays
Using the formula of the gravitational wave energy spectrum shown in [] and employing the analysis methods used in [], we give the constraint on the tensor-to-scalar ratio ${\rm{log}_{10}}r=0.485^{+0.013}_{-0.010}$ at the $1\,\sigma$ CL, ${\rm{log}_{10}}r=0.485^{+0.023}_{-0.028}$ at the $2\,\sigma$ CL and ${\rm{log}_{10}}r=0.485^{+0.027}_{-0.031}$ at the $3\,\sigma$ CL via the international pulsar timing array data release two (IPTA DR2) []. One can easily find that they are far larger than the CMB constraints and consequently, PTA is in a very strong tension with CMB when constraining the single-field slow-roll inflation.

### v2

@@ Abstract
Since our constraining results have revealed that the GUT scale can be reached during inflation within $3\,\sigma$ level, the particle production of dark matter around the GUT scale could be allowed by observations. As a consequence, our constraints could inspire the exploration of dark matter model building at extremely high energy scales, and intriguingly, provide a new window to investigate the physics before inflation.
@@
Acknowledgements. DW is supported by the CDEIGENT Fellowship of Consejo Superior de Investigaciones Científicas (CSIC). DW warmly thanks the usage of Planck Planck:2018vyg , ACT ACT:2020gnv , SPT SPT-3G:2022hvq , BK18 BICEP:2021xfz and DESI DESI:2024mwx data.